HUBBLE SPACE TELESCOPE DAILY REPORT #2983 PERIOD COVERED: 0000Z (UTC) 10/23/01 - 0000Z (UTC) 10/24/01 Daily Status Report as of 297/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Four Sets of STIS/CCD/MA1/MA2 9116 (Understanding High-Redshift and Starburst Galaxies: A UV Spectroscopic Survey of B- Stars in the SMC) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was used to construct a high quality spectral library at low metallicity in order to synthesize the UV spectra of high redshift star-forming galaxies and nearby starbursts. No problems were reported. 1.2 Completed WF/PC-2 9249 (POMS Test Proposal: WFII Backup Parallel Archive Proposal II) The WF/PC-2 was used to execute a POMS test proposal designed to simulate scientific plans. The proposal completed with no reported problems. 1.3 Completed Three Sets of WF/PC-2 8059 (POMS Test Proposal: Targeted Parallel Archive Proposal) The WF/PC-2 was used to observe the parallel opportunities available in the neighborhood of bright galaxies are treated in a slightly different way from the normal pure parallels. Local Group galaxies offer the opportunity for a closer look at young stellar populations. Narrow-band images in F656N can be used both to identify young stars via their emission lines, and to map the gas distribution in star-forming regions. The observations completed nominally. 1.4 Completed Two Sets of STIS/CCD 8901 (Dark Monitor-Part 1) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the darks. There was no anomalous activity. 1.5 Completed STIS/CCD 8903 (Bias Monitor - Part 1) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4 to build up high-S/N superbiases and track the evolution of hot columns. There were no problems. 1.6 Completed STIS/CCD 8262 (Nuclear Kinematics of the Dense Globular Cluster M-15) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of star NGC7078-CE as part of the study of the nuclear kinematics of the dense globular cluster M-15. The high central density of M-15 makes it unique for probing the structure and evolution of globular clusters. The observations were completed as planned, and no anomalies were reported. 1.7 Completed Eight Sets of WF/PC-2 9244 (POMS Test Proposal: WFII Parallel Archive Proposal Continuation) The WF/PC-2 was used to perform a generic target version of the Archival Pure Parallel program. The program will be used to take parallel images of random areas of the sky, following the recommendations of the Parallels Working Group. The observations completed with no anomalous activity. 1.8 Completed STIS/CCD 9143 (Spectrophotometry of Nearby Seyfert 2 Nuclei: Can We Eliminate the Seyfert 2 Class?) The Space Telescope Imaging Spectrograph (CCD) was used to investigate Seyfert 2s that are distinguished by the absence of the broad emission lines characteristic of Seyfert 1s and more luminous QSOs. Are Seyfert 2s fundamentally different from Seyfert 1s and their brighter cousins? Or is the broad emission line region in Seyfert 2s simply suppressed by obscuring material as postulated by the unification model? If the latter model is correct, the weak broad emission lines in the Seyfert 2s may simply be overwhelmed by starlight from the circumnuclear region, particularly in the case of recent star formation. It is proposed to determine if all Seyfert 2s have {weak} broad emission line regions by obtaining long-slit STIS spectroscopy for a well-defined sample of 20 Seyfert 2s {3 archival, 17 new}. The observations completed with no anomalous activity. 1.9 Completed STIS/CCD 9088 (Next Generation Spectral Library of Stars) The Space Telescope Imaging Spectrograph (CCD) was used to produce a "Next Generation'' Spectral Library of 600 stars for use in modeling the integrated light of galaxies and clusters by using the low dispersion UV and optical gratings of STIS. The library will be roughly equally divided among four metallicities, very low {Fe/H < -1.5}, low {-1.5 < Fe/H < -0.5}, near-solar {-0.5 < Fe/H < 0.1}, and super-solar {Fe/H > 0.1}, well-sampling the entire HR-diagram in each bin. Such a library will surpass all extant compilations and have lasting archival value, well into the Next Generation Space Telescope era. No problems occurred. 1.10 Completed WF/PC-2 8941 (Cycle 10 UV Earthflats) The WF/PC-2 was used to monitor flat field stability by obtaining sequences of earth streak flats to improve the quality of pipeline flat fields for the WFPC2 UV filter set. The proposal had no problems. 1.11 Completed STIS/CCD 9077 (Survey of the LMC Planetary Nebulae) The Space telescope Imaging Spectrograph (CCD) was used to perform a snapshot survey of all known LMC planetary nebulae {PNe} in order to study the co-evolution of the nebulae and their central stars, and to probe the chemical enrichment history of the LMC. There were no reported problems. 1.12 Completed STIS/CCD 7912 (STIS Parallel Archive Proposal - Nearby Galaxies - Imaging and Spectroscopy) The Space Telescope Imaging Spectrograph (CCD) was used to make parallel observations of nearby galaxies. This survey will be useful to study the star formation histories, chemical evolution, and distances to these galaxies. These data will be placed immediately into the Hubble Data Archive. The observations were completed as scheduled, and no problems were reported. 1.13 Completed WF/PC-2 8936 (Cycle 10 Supplemental Darks Pt1/3) The WF/PC-2 was used to perform a dark calibration program that obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The proposal completed with no reported problems. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 9 Successful: 9 Scheduled Re-acquisitions: 3 Successful: 3 2.2 FHST Updates: Scheduled: 16 Successful: 16 2.3 Operations Notes: The SSR EDAC error counter was cleared twice per ROP SR-1A. There was a STIS EMC re-try at 296/171713Z. Using ROP NS-12, the STIS flight software error counter was reset at 296/1721Z. The engineering status buffer limits were adjusted at 296/2222Z per ROP DFD-18A. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /CAW